# Using Anomaly Detection to Detect Poisoning Attacks in Federated Learning Applications

### v3

@@ Abstract
Adversarial attacks such as poisoning attacks have attracted the attention of many machine learning researchers. Traditionally, poisoning attacks attempt to inject adversarial training data in order to manipulate the trained model. In federated learning (FL), data poisoning attacks can be generalized to model poisoning attacks, which cannot be detected by simpler methods due to the lack of access to local training data by the detector. State-of-the-art poisoning attack detection methods for FL have various weaknesses, e.g., the number of attackers has to be known or not high enough, working with i.i.d. data only, and high computational complexity. To overcome above weaknesses, we propose a novel framework for detecting poisoning attacks in FL, which employs a reference model based on a public dataset and an auditor model to detect malicious updates. We implemented a detector based on the proposed framework and using a one-class support vector machine (OC-SVM), which reaches the lowest possible computational complexity $\mathcal{O}(K)$ where $K$ is the number of clients. We evaluated our detector’s performance against state-of-the-art (SOTA) poisoning attacks for two typical applications of FL: electrocardiograph (ECG) classification and human activity recognition (HAR). Our experimental results validated the performance of our detector over other SOTA detection methods.
Index Terms:
Federated learning, security, privacy, anomaly detection, poisoning attacks, data poisoning, model poisoning, Byzantine attacks.
@@ 1 Introduction
Privacy and security are among the top issues to be addressed in privacy and security-sensitive applications of machine learning (ML), such as healthcare, autonomous vehicles, etc. Federated learning (FL) has been introduced to enhance data privacy in machine learning applications [1]. FL attempts to provide enhanced privacy to the data owners by collaboratively training a joint model by only sharing parameters of locally trained models, in this way the data owners never share raw data, but can collaboratively train a joint robust global model. For instance, various hospitals can train ML models jointly for healthcare applications without directly sharing their privacy-sensitive data with each other. Generally speaking, FL iterates in three steps: the global server (i.e., a cloud server), which maintains the global model, sends the global model to the edge devices; the edge devices update the local models using their local training data and share the trained parameters of the locally trained model with the global server, and the global server updates the global model by incorporating the shared parameters according to an aggregation rule. For example, the mean aggregation rule, which computes the weighted average of the shared local model’s parameters, is one of the widely used aggregation algorithms [2]. Nevertheless, in such cases, the global model can be easily manipulated, even if a single-edge device is compromised [3, 4, 5]. The attack surface of FL is growing due to its distributed nature. For example, malicious peers can launch data poisoning [6, 7] or model poisoning [8] attacks, in which one or more malicious edge devices manipulate their local training data or the local model trained on benign data, to impair the performance of the updated global model.
@@
FL can be divided into three phases: data and behavior auditing, training, and testing. FL faces different kinds of security threats in each phase [9]. Hence, establishing secure FL needs to take effective measures at each phase to mitigate such threats. A solution before integrating a local model into the global model is to audit the local data before training the local model(s). However, due to privacy concerns and the architecture of FL, it is challenging to conduct such audits [9]. A trivial method to address model poisoning attacks could be using accuracy, i.e., using accuracy as a measure to examine the quality of data being used to train the local model. Nevertheless, such methods cannot be generalized as accuracy alone cannot reveal information about the underlying data. Just looking at the accuracy it is hard to claim that a local model is trained on benign or malicious data. Furthermore, models can be designed to have high accuracy for the testing samples by including them in the training dataset. A model can have low accuracy even if it has been trained on benign data depending on the amount of training data, training epochs, hyper-parameters tuning, optimization, etc. Hence, new solutions are required to detect such model and data poisoning attacks. Methods should be developed to verify that the shared local model gradients are not trained on anomalous or malicious (e.g., featured poisoning, label poisoning) data. In other words, malicious behaviors of local models should be detected before considering them in the global aggregation process in order to prevent malicious peers from compromising and manipulating the global model. Therefore, many researchers have proposed different poisoning attack detection methods for FL [10, 11, 12], which unfortunately all suffer from various weaknesses, e.g., the number of attackers has to be known or cannot be too high, working with i.i.d. data only, and high computational complexity. To overcome the above weaknesses of the state-of-the-art (SOTA) methods, our work made the following key contributions:
@@
The rest of the paper is organized as follows. The next section introduces the background and related work. Section 3 presents the proposed framework. Sections 4 and 5 cover the performance evaluation and comparison of the proposed framework, respectively. Section 6 presents some further discussions and discuss limitations of our work. Finally, the last section concludes the paper.
@@ 2.1 Federated Learning
Federated learning (FL) [1] is a concept of distributed machine learning in which different edge devices or silos (hospitals, companies, etc.) collaborate to train a joint model known as a global model. The global model is trained without requiring participating edge devices to share their local raw data with other participating nodes or a centralized server. Instead, each edge device trains their own local model using their local data and shares only the trained parameters with a central server in centralized FL or with all other nodes in decentralized FL. The aggregation of local models are done according to a given aggregation algorithm to compute parameters of the global model [13], among which FedAvg [14] is the most commonly used algorithm [2]. Mathematically, FedAvg is given by the following equation:
@@ 2.2 Stitching Connectivity
Stitching connectivity [15] is a method to measure the similarity of internal representations of different models trained using different but similar data. Consider two models A and B, which have the same architecture. For A and B to be stitching connected, they can be stitched at all the layers to each other. In other words, two models, let’s say A and B with identical architecture but trained using stochastic gradient descent(or its variants) using independent random seeds and independent training sets taken from the same distribution. Then the two trained models are stitched and connected for natural architectures and data distributions. Hence, we expect the models trained on similar but different training sets of the same distribution will behave similarly. Based on this property, we adopt Proposition 1.
@@ 2.4 Byzantine attacks
Byzantine attacks are a type of attacks where a trusted device or a set of devices turn rogue and try to compromise the overall system. Such attacks can significantly reduce the performance of the global model in federated learning [17]. Researchers have shown that in the case of some aggregation algorithms even the presence of a single malicious node can significantly reduce the performance of the global model [3, 4, 5]. Byzantine attacks, such as poisoning attacks [18, 19], can substantially reduce the performance (classification accuracy, precision, and recall) of FedAvg, even in the presence of a very small percentage of adversarial participants in the network. Such attacks can be classified as targeted attacks that negatively impact only one or more target (but not all) classes under attack, and untargeted attacks that impact all the classes negatively. Furthermore, poisoning attacks are mainly classified into two main categories: data poisoning [18] and model poisoning [3] attacks depending on the phase where the attacks are executed. If the attacker manipulates the training data then this is called data poisoning attacks and if the attacker manipulates the trained model’s parameters then such attacks are called model poisoning attacks. Further details of each type of poisoning are given as follows:
@@
Data poisoning attacks [18, 17, 20] are those attacks in which the attacker manipulates the training data directly according to a given strategy and then trains the model using the manipulated dataset. In this study we consider the following four types of SOTA data poisoning attacks:
@@
    Gradient ascent attacks: In such attacks the attacker trains the models using a gradient ascent instead of gradient descent optimizer.
@@ 2.5 Existing defense mechanisms
In order to address the above-mentioned attacks, many defense methods have been proposed. Hu et al. [17] conducted a survey of the state-of-the-art defensive methods against byzantine attacks in FL. They show that byzantine attacks by malicious clients can significantly reduce the accuracy of the global model. Moreover, their results show that existing defense solutions cannot fully be protecting FL against such attacks. In [12] Xia et al. summarized poisoning attacks and defense strategies according to their methods and targets. In this section, we discuss some of the existing SOTA defense mechanisms below.
@@ 2.5.1 Distance-Based Mechanisms
Such mechanisms detect malicious updates by calculating the distance between updates. Updates with a larger distance from others are discarded from global aggregation. For example, Blanchard et al. [4] proposed Krum, where the central server selects updates with minimum distance from the neighbors. Similarly, Xia et al. [21] proposed a method to discard the shared parameters with a large distance from the mean of shared parameters. Bo et al. [22] proposed FedInv, which inverses the updates from each client to generate a dummy dataset. The server then calculates Wasserstein distances for each client and removes the update(s) with exceptional Wasserstein distances from others. Gupta et al. [23] proposed a method called MUD-HoG, a method to address poisoning attacks in federated learning using long-short history of gradients of clients. Nevertheless, for such methods to work the number of attackers is needed to be known in advance or needs to be less than a certain percentage of total clients in the network.
@@ 2.5.3 Statistical Mechanisms
A method in this category use statistical features of the shared gradients or updates. Commonly used features are mean or median. Such features can circumvent benign updates and help to achieve a robust and probably more benign gradient. For example, a coordinate-wise median and a coordinate-wise median-based solution are provided by Yin et al. [26], which aggregates the parameters of local models independently, i.e., for the $i$-th model parameter the global server sorts the $i$-th parameter of the $m$ other local models. Removes the largest and smallest parameters and computes the mean of the remaining parameters as the $i$-th parameter of the global model. Similarly, El Mhamdi et al. [27] proposed to combine Krum and a variant of trimmed mean [26]. However, these approaches are vulnerable to poisoning attacks even using robust aggregation [9]. Additionally, for such methods to work, the number of attackers should be limited by an upper bound (e.g., 50%) or should be known prior.
@@ 2.5.4 Target Optimization-Based Mechanisms
Target optimization-based methods optimize an objective function to improve the robustness of the global model. For example, Li et al. [28] RSA which regularizes the objective in such a way that it forces each local model in federated learning to be close to the global model. However, such methods only address data poisoning attacks and fail to eliminate model poisoning attacks.
@@ 3.1 Threat Model
    There are one or more malicious edge devices who try to launch a model poisoning attack against the global model, possibly in a collaborative manner (i.e., launching a colluding attack).
@@
    All the attackers have two strategies: manipulating their local training data and train the local model using the manipulated data, and manipulating the model parameters after training it on benign data.
@@ 3.2 Overview
In this subsection, we describe the proposed framework. An overview of the proposed method has been shown in Figure 1. Let us assume $K$ edged devices (hospitals, organizations, etc.) collaborate to train a joint global model GM. An edge $\textbf{E}_{k}$ trains a local model $\textbf{LM}_{k}$ using its local data $\textbf{D}_{k}$, where $k={1,2,\dots,K}$. The global server GS is responsible for receiving the updates from edge devices and aggregation. We assume that a trusted third party or GS (We consider trusted third party to be a different entity from GS) also has an open-source dataset which is called public data and we represent it as DP. DP is supposed to be a representative dataset of all the classes in a classification problem, i.e., it has samples from each candidate class. The training for a global round is given as follows.
@@
    After training, trusted third party makes predictions with RM using $\textbf{DP}_{\text{train}}$ and $\textbf{DP}_{\text{test}}$. During the predictions for each dataset, trusted third party taps the activations of the last hidden layer for each input sample using Algorithm 1 to create a dataset $\textbf{DA}_{\text{train}}$, and $\textbf{DA}_{\text{test}}$ for each $\textbf{DP}_{\text{train}}$ and $\textbf{DP}_{\text{test}}$, respectively.
@@ 10       Append $s$ in $\textbf{DA}_{i}$ as a new data sample
Based on Proposition 2, we expect that the models trained on different datasets of similar distributions will behave similarly. Since the information specific to the data is learned by the higher-level layers of the model, we take the activation of the last hidden layer (in the case of CNN’s last convolutional layer) to capture more information (features) related to the training data. We only consider the activations of the last as it provides sufficient information about the underlying training data to detect poisoning attacks. Activations from other higher-level layers can also be incorporated for this purpose, nevertheless, this increases the computational costs with no significant improvement in the results for detecting poisoning attacks. Hence, we only consider the activations of the last convolutional hidden layer. We capture this property by training another model (AM) to learn the behavior of a model trained on the benign dataset (RM in our case). Hence, models behaving similarly to the model trained on the benign datasets are probably free from poisoning attacks, and at least they do not degrade the performance of the global model.
@@ 4.1 Experimental Setup
Classifiers: We developed a convolution neural networks (CNN) based classifier for each application. The developed classifiers do not achieve the optimum classification for the considered datasets. This is because our objective is to show that our proposed framework can detect anomalous (poisoning attacks), not to achieve the best performance in terms of classification. For ECG classification we developed a five-class classifier and the aim of FL here is to learn a global five-class classifier, and for HAR we developed a fourteen-class classifier, and the aim of FL here is to learn a global fourteen-class classifier.
@@ 4.2 Performance Evaluation
In order to evaluate the performance of the proposed framework, we first tested the proposed framework to check its ability to differentiate samples in $\textbf{DA}_{i}$ of benign edges from samples in $\textbf{DA}_{i}$ of malicious edges. We trained a RM and an AM using a public dataset (which is excluded from the training data of the edge devices and the test data). We distributed each candidate dataset (ECG and HAR) equally (in the case of the HAR dataset, benign nodes contain slightly more data) but randomly among the benign and malicious nodes. For each dataset, we did the following. We created $\textbf{DA}_{i}$ using each benign edge and labeled each sample as 1 (for benign). Similarly, for the malicious edge, we created $\textbf{DA}_{i}$ using each type of model and data poisoning attack discussed in Section 2.4 and selected random samples from each attack’s $\textbf{DA}_{i}$ to make a new $\textbf{DA}_{i}$ and labeled each sample as -1. We combined the two $\textbf{DA}_{i}$’s that we created. Then, this lastly created dataset was used to test the AM. Note that here we just show the ability of the proposed framework to classify benign and malicious samples not the detection of updates. We will show the detection of updates later. Table I shows the classification accuracy of the proposed framework to differentiate samples generated using shared parameters of benign edged devices and the samples generated using shared parameters of malicious edge devices for ECG classification. Similarly, Table II shows the classification accuracy of the proposed framework to differentiate samples generated using shared parameters of benign edged devices and the samples generated using shared parameters of malicious edge devices for HAR. For both types of applications, it can be seen that the proposed framework can differentiate samples of benign and malicious edged devices very well, with an overall accuracy of 94% and 99% for HAR and ECG classification, respectively. As mentioned previously, here we calculate the overall accuracy of AM to classify benign and malicious samples which is 94% and 99% for HAR and ECG respectively. It should be noted that the poison attack detection is still 100% as updates with 94% and 99% poisoned samples will have a poison rate of around 94% and 99%, respectively. Hence, updates with 94% and 99% poison rates will be classified as malicious and will be removed from global aggregation.
TABLE I: Classification accuracy of AM for ECG classification
@@
TABLE II: Classification accuracy of AM for HAR
@@
Now, we show the ability of the proposed framework to detect malicious updates and remove them from the global aggregation. We simulated the proposed framework as follows. First, the trusted third party prepares the RM and the AM and trains the RM and the AM using the public dataand sends it to GS. Then the benign devices and the attacker send the updates to the GS. The attacker launches a given attack at each global round. For example, for round $r$ it uses a random label flipping attack and for $r+1$ the attacker launches a feature poisoning attack, and so on. We tested the performance of our proposed framework under all the different types of model and data poisoning attacks listed in Section 2.4. Moreover, we distributed the data among the edged devices randomly but using non-independent and identically distributed unbalanced data distribution, where each edge has 40-50% fewer data for a given class. For example, edge 1 has 40% fewer samples of non-ecotic beats class for the ECG dataset as compared to other edge devices, and edge 2 has 50% fewer samples of supraventricular ectopic beats class compared to other edge devices. In comparison, edge 3 (attacker) has an equal number of samples for each class. This is done in order to give more attacking power to the edge3. For example, if the attacker has an equal number of samples of each class it can launch a random label flipping attack which will eventually affect all the classes and will affect the global model significantly. We follow similar distribution for the HAR dataset.
@@
First, we present the comparison of the accuracy of the global model after one global round with and without our proposed framework, under four different data poisoning attacks. Note that our global model achieves optimum accuracy after one global round, hence we launched the attack in each global round. Otherwise, the attacks can be launched at any given global round. Figure 2 presents a comparison of the accuracy of the global model under different data poisoning attacks and with and without the proposed framework for ECG classification. Here, GM is a global model with our proposed framework under different data poisoning attacks, LS represents the global model under a label swapping attack without our proposed framework, RLF represents the global model under a random label and feature poisoning attack without our proposed framework, RL represents the global model under a random label attack without our proposed framework, and FP represents the global model under a feature poisoning attack without our proposed framework, all in a federated setting. It can be seen that the performance of the global model deteriorates significantly under the data poisoning attacks. It can also be seen that the performance of the global model was not affected by the poisoning attacks when we applied our proposed framework that could detect and eliminate almost all poisoned updates from the malicious edge before global aggregation. As mentioned previously, even if the accuracy of AM for benign and malicious sample detection is not 100%, it can accurately detect all the malicious updates because the poisoning rate is high. Hence, its malicious updates detection is still 100%. Similarly, Figure 3 presents the performance of the global model compared under four different data poisoning attacks with and without adopting the proposed framework for HAR. Since all the malicious updates are detected and removed before aggregation, the accuracy of the global model remains almost similar for a given dataset.
@@
Moreover, Figure 4 presents a comparison of the accuracy of the global model under four different model poisoning attacks with and without our proposed framework for ECG classification. Here, GM shows the performance of the global model with our proposed framework, and SF shows the performance of the global model under a sign flip attack without our proposed framework, SV shows the performance of the global model under a same value attack without our proposed framework, AGA shows the performance of the global model under a performance additive Gaussian noise attack without a proposed framework, and GA shows the performance of the global model under a gradient ascent attack without our proposed framework, all in a federated setting. Similarly, Figure 5 presents a comparison of the accuracy of the global model under the four different model poisoning attacks with and without our proposed framework for HAR.

### v4

@@ Abstract
Adversarial attacks such as poisoning attacks have attracted the attention of many machine learning researchers. Traditionally, poisoning attacks attempt to inject adversarial training data to manipulate the trained model. In federated learning (FL), data poisoning attacks can be generalized to model poisoning attacks, which cannot be detected by simpler methods due to the lack of access to local training data by the detector. State-of-the-art poisoning attack detection methods for FL have various weaknesses, e.g., the number of attackers has to be known or not high enough, working with i.i.d. data only, and high computational complexity. To overcome the above weaknesses, we propose a novel framework for detecting poisoning attacks in FL, which employs a reference model based on a public dataset and an auditor model to detect malicious updates. We implemented a detector based on the proposed framework and using a one-class support vector machine (OC-SVM), which reaches the lowest possible computational complexity $\mathcal{O}(K)$ where $K$ is the number of clients. We evaluated our detector’s performance against state-of-the-art (SOTA) poisoning attacks for two typical applications of FL: electrocardiograph (ECG) classification and human activity recognition (HAR). Our experimental results validated the performance of our detector over other SOTA detection methods.
keywords:
[…]
@@ 1 Introduction
Privacy and security are among the top issues to be addressed in privacy and security-sensitive applications of machine learning (ML), such as healthcare, autonomous vehicles, etc. Federated learning (FL) has been introduced to enhance data privacy in machine learning applications [1]. FL attempts to provide enhanced privacy to the data owners by collaboratively training a joint model by only sharing parameters of locally trained models, in this way the data owners never share raw data, but can collaboratively train a joint robust global model. For instance, various hospitals can train ML models jointly for healthcare applications without directly sharing their privacy-sensitive data. Generally speaking, FL iterates in three steps: the global server (i.e., a cloud server), which maintains the global model, sends the global model to the edge devices; the edge devices update the local models using their local training data and share the trained parameters of the locally trained model with the global server, and the global server updates the global model by incorporating the shared parameters according to an aggregation rule. For example, the mean aggregation rule, which computes the weighted average of the shared local model’s parameters, is one of the widely used aggregation algorithms [2]. Nevertheless, in such cases, the global model can be easily manipulated, even if a single-edge device is compromised [3, 4, 5]. The attack surface of FL is growing due to its distributed nature. For example, malicious peers can launch data poisoning [6, 7] or model poisoning [8] attacks, in which one or more malicious edge devices manipulate their local training data or the local model trained on benign data, to impair the performance of the updated global model.
@@
FL can be divided into three phases: data and behavior auditing, training, and testing. FL faces different kinds of security threats in each phase [9]. Hence, establishing secure FL needs to take effective measures at each phase to mitigate such threats. A solution before integrating a local model into the global model is to audit the local data before training the local model(s). However, due to privacy concerns and the architecture of FL, it is challenging to conduct such audits [9]. A trivial method to address model poisoning attacks could be using accuracy, i.e., using accuracy as a measure to examine the quality of data being used to train the local model. Nevertheless, such methods cannot be generalized as accuracy alone cannot reveal information about the underlying data. Just looking at the accuracy it is hard to claim that a local model is trained on benign or malicious data. Furthermore, models can be designed to have high accuracy for the testing samples by including them in the training dataset. A model can have low accuracy even if it has been trained on benign data depending on the amount of training data, training epochs, hyper-parameters tuning, optimization, etc. Hence, new solutions are required to detect such model and data poisoning attacks. Methods should be developed to verify that the shared local model gradients are not trained on anomalous or malicious (e.g., featured poisoning, label poisoning) data. In other words, malicious behaviors of local models should be detected before considering them in the global aggregation process to prevent malicious peers from compromising and manipulating the global model. Therefore, many researchers have proposed different poisoning attack detection methods for FL [10, 11, 12], which unfortunately all suffer from various weaknesses, e.g., the number of attackers has to be known or cannot be too high, working with i.i.d. data only, and high computational complexity. To overcome the above weaknesses of the state-of-the-art (SOTA) methods, our work made the following key contributions:
@@
The rest of the paper is organized as follows. The next section introduces the background and related work. Section 3 presents the proposed framework. Sections 4 and 5 cover the performance evaluation and comparison of the proposed framework, respectively. Section 6 presents some further discussions and discusses the limitations of our work. Finally, the last section concludes the paper.
@@ 2.1 Federated Learning
Federated learning (FL) [1] is a concept of distributed machine learning in which different edge devices or silos (hospitals, companies, etc.) collaborate to train a joint model known as a global model. The global model is trained without requiring participating edge devices to share their local raw data with other participating nodes or a centralized server. Instead, each edge device trains its local model using its local data and shares only the trained parameters with a central server in centralized FL or with all other nodes in decentralized FL. The aggregation of local models is done according to a given aggregation algorithm to compute parameters of the global model [13], among which FedAvg [14] is the most commonly used algorithm [2]. Mathematically, FedAvg is given by the following equation:
@@ 2.2 Stitching Connectivity
Stitching connectivity [15] is a method to measure the similarity of internal representations of different models trained using different but similar data. Consider two models A and B, which have the same architecture. For A and B to be stitched connected, they can be stitched at all the layers to each other. In other words, two models, let’s say A and B with identical architecture but trained using stochastic gradient descent(or its variants) using independent random seeds and independent training sets taken from the same distribution. Then the two trained models are stitched and connected for natural architectures and data distributions. Hence, we expect the models trained on similar but different training sets of the same distribution will behave similarly. Based on this property, we adopt Proposition 1.
@@ 2.4 Byzantine attacks
Byzantine attacks are a type of attack where a trusted device or a set of devices turn rogue and try to compromise the overall system. Such attacks can significantly reduce the performance of the global model in federated learning [17]. Researchers have shown that in the case of some aggregation algorithms, even the presence of a single malicious node can significantly reduce the performance of the global model [3, 4, 5]. Byzantine attacks, such as poisoning attacks [18, 19], can substantially reduce the performance (classification accuracy, precision, and recall) of FedAvg, even in the presence of a very small percentage of adversarial participants in the network. Such attacks can be classified as targeted attacks that negatively impact only one or more target (but not all) classes under attack, and untargeted attacks that impact all the classes negatively. Furthermore, poisoning attacks are mainly classified into two main categories: data poisoning [18] and model poisoning [3] attacks depending on the phase where the attacks are executed. If the attacker manipulates the training data then this is called data poisoning attacks and if the attacker manipulates the trained model’s parameters then such attacks are called model poisoning attacks. Further details of each type of poisoning are given as follows:
@@
Data poisoning attacks [18, 17, 20] are those attacks in which the attacker manipulates the training data directly according to a given strategy and then trains the model using the manipulated dataset. In this study, we consider the following four types of SOTA data poisoning attacks:
@@
    Gradient ascent attacks: In such attacks the attacker trains the models using a gradient ascent instead of a gradient descent optimizer.
@@ 2.5 Existing defense mechanisms
To address the above-mentioned attacks, many defense methods have been proposed. Hu et al. [17] surveyed the state-of-the-art defensive methods against byzantine attacks in FL. They show that byzantine attacks by malicious clients can significantly reduce the accuracy of the global model. Moreover, their results show that existing defense solutions cannot fully protect FL against such attacks. In [12] Xia et al. summarized poisoning attacks and defense strategies according to their methods and targets. In this section, we discuss some of the existing SOTA defense mechanisms below.
@@ 2.5.1 Distance-Based Mechanisms
Such mechanisms detect malicious updates by calculating the distance between updates. Updates with a larger distance from others are discarded from global aggregation. For example, Blanchard et al. [4] proposed Krum, where the central server selects updates with minimum distance from the neighbors. Similarly, Xia et al. [21] proposed a method to discard the shared parameters with a large distance from the mean of shared parameters. Bo et al. [22] proposed FedInv, which inverses the updates from each client to generate a dummy dataset. The server then calculates Wasserstein distances for each client and removes the update(s) with exceptional Wasserstein distances from others. Gupta et al. [23] proposed a method called MUD-HoG, a method to address poisoning attacks in federated learning using a long-short history of gradients of clients. Nevertheless, for such methods to work the number of attackers is needed to be known in advance or needs to be less than a certain percentage of total clients in the network.
@@ 2.5.3 Statistical Mechanisms
A method in this category uses statistical features of the shared gradients or updates. Commonly used features are mean or median. Such features can circumvent benign updates and help to achieve a robust and probably more benign gradient. For example, a coordinate-wise median and a coordinate-wise median-based solution are provided by Yin et al. [26], which aggregates the parameters of local models independently, i.e., for the $i$-th model parameter the global server sorts the $i$-th parameter of the $m$ other local models. Removes the largest and smallest parameters and computes the mean of the remaining parameters as the $i$-th parameter of the global model. Similarly, El Mhamdi et al. [27] proposed to combine Krum and a variant of trimmed mean [26]. However, these approaches are vulnerable to poisoning attacks even using robust aggregation [9]. Additionally, for such methods to work, the number of attackers should be limited by an upper bound (e.g., 50%) or should be known prior.
@@ 2.5.4 Target Optimization-Based Mechanisms
Target optimization-based methods optimize an objective function to improve the robustness of the global model. For example, Li et al. [28] RSA, which regularizes the objective in such a way that it forces each local model in federated learning to be close to the global model. However, such methods only address data poisoning attacks and fail to eliminate model poisoning attacks.
@@ 3.1 Threat Model
    There are one or more malicious edge devices that try to launch a model poisoning attack against the global model, possibly in a collaborative manner (i.e., launching a colluding attack).
@@
    All the attackers have two strategies: manipulating their local training data and training the local model using the manipulated data, and manipulating the model parameters after training it on benign data.
@@ 3.2 Overview
In this subsection, we describe the proposed framework. An overview of the proposed method has been shown in Figure 1. Let us assume $K$ edged devices (hospitals, organizations, etc.) collaborate to train a joint global model GM. An edge $\textbf{E}_{k}$ trains a local model $\textbf{LM}_{k}$ using its local data $\textbf{D}_{k}$, where $k={1,2,\dots,K}$. The global server GS is responsible for receiving the updates from edge devices and aggregation. We assume that a trusted third party or GS (We consider a trusted third party to be a different entity from GS) also has an open-source dataset which is called public data and we represent it as DP. DP is supposed to be a representative dataset of all the classes in a classification problem, i.e., it has samples from each candidate class. The training for a global round is given as follows.
@@
    After training, trusted third party makes predictions with RM using $\textbf{DP}_{\text{train}}$ and $\textbf{DP}_{\text{test}}$. During the predictions for each dataset, a trusted third party taps the activations of the last hidden layer for each input sample using Algorithm 1 to create a dataset $\textbf{DA}_{\text{train}}$, and $\textbf{DA}_{\text{test}}$ for each $\textbf{DP}_{\text{train}}$ and $\textbf{DP}_{\text{test}}$, respectively.
@@ 10       Append $s$ in $\textbf{DA}_{i}$ as a new data sample
Based on Proposition 2, we expect that the models trained on different datasets of similar distributions will behave similarly. Since the higher-level layers learn the information specific to the data of the model, we take the activation of the last hidden layer (in the case of CNN’s last convolutional layer) to capture more information (features) related to the training data. We only consider the activations of the last as it provides sufficient information about the underlying training data to detect poisoning attacks. Activations from other higher-level layers can also be incorporated for this purpose, nevertheless, this increases the computational costs with no significant improvement in the results for detecting poisoning attacks. Hence, we only consider the activations of the last convolutional hidden layer. We capture this property by training another model (AM) to learn the behavior of a model trained on the benign dataset (RM in our case). Hence, models behaving similarly to the model trained on the benign datasets are probably free from poisoning attacks, and at least they do not degrade the performance of the global model.
@@ 4.1 Experimental Setup
Classifiers: We developed a convolution neural networks (CNN) based classifier for each application. The developed classifiers do not achieve the optimum classification for the considered datasets. This is because our objective is to show that our proposed framework can detect anomalous (poisoning attacks), not to achieve the best performance in terms of classification. For ECG classification, we developed a five-class classifier, and the aim of FL here is to learn a global five-class classifier. for HAR, we developed a fourteen-class classifier, and the aim of FL here is to learn a global fourteen-class classifier.
@@ 4.2 Performance Evaluation
To evaluate the performance of the proposed framework, we first tested the proposed framework to check its ability to differentiate samples in $\textbf{DA}_{i}$ of benign edges from samples in $\textbf{DA}_{i}$ of malicious edges. We trained a RM and an AM using a public dataset (which is excluded from the training data of the edge devices and the test data). We distributed each candidate dataset (ECG and HAR) equally (in the case of the HAR dataset, benign nodes contain slightly more data) but randomly among the benign and malicious nodes. For each dataset, we did the following. We created $\textbf{DA}_{i}$ using each benign edge and labeled each sample as 1 (for benign). Similarly, for the malicious edge, we created $\textbf{DA}_{i}$ using each type of model and data poisoning attack discussed in Section 2.4 and selected random samples from each attack’s $\textbf{DA}_{i}$ to make a new $\textbf{DA}_{i}$ and labeled each sample as -1. We combined the two $\textbf{DA}_{i}$’s that we created. Then, this lastly created dataset was used to test the AM. Note that here we just show the ability of the proposed framework to classify benign and malicious samples not the detection of updates. We will show the detection of updates later. Table 1 shows the classification accuracy of the proposed framework to differentiate samples generated using shared parameters of benign edged devices and the samples generated using shared parameters of malicious edge devices for ECG classification. Similarly, Table 2 shows the classification accuracy of the proposed framework to differentiate samples generated using shared parameters of benign edged devices and the samples generated using shared parameters of malicious edge devices for HAR. For both types of applications, it can be seen that the proposed framework can differentiate samples of benign and malicious edged devices very well, with an overall accuracy of 94% and 99% for HAR and ECG classification, respectively. As mentioned previously, here we calculate the overall accuracy of AM to classify benign and malicious samples which is 94% and 99% for HAR and ECG respectively. It should be noted that the poison attack detection is still 100% as updates with 94% and 99% poisoned samples will have a poison rate of around 94% and 99%, respectively. Hence, updates with 94% and 99% poison rates will be classified as malicious and will be removed from global aggregation.
Table 1: Classification accuracy of AM for ECG classification
@@
Table 2: Classification accuracy of AM for HAR
@@
Now, we show the ability of the proposed framework to detect malicious updates and remove them from the global aggregation. We simulated the proposed framework as follows. First, the trusted third party prepares the RM and the AM and trains the RM and the AM using the public data and sends it to GS. Then the benign devices and the attacker send the updates to the GS. The attacker launches a given attack at each global round. For example, for round $r$ it uses a random label flipping attack and for $r+1$ the attacker launches a feature poisoning attack, and so on. We tested the performance of our proposed framework under all the different types of model and data poisoning attacks listed in Section 2.4. Moreover, we distributed the data among the edged devices randomly but using non-independent and identically distributed unbalanced data distribution, where each edge has 40-50% fewer data for a given class. For example, edge 1 has 40% fewer samples of non-ecotic beats class for the ECG dataset as compared to other edge devices, and edge 2 has 50% fewer samples of supraventricular ectopic beats class compared to other edge devices. In comparison, edge 3 (attacker) has an equal number of samples for each class. This is done to give more attacking power to the edge3. For example, if the attacker has an equal number of samples of each class it can launch a random label-flipping attack which will eventually affect all the classes and will affect the global model significantly. We follow a similar distribution for the HAR dataset.
@@
First, we present the comparison of the accuracy of the global model after one global round with and without our proposed framework under four different data poisoning attacks. Note that our global model achieves optimum accuracy after one global round, hence we launched the attack in each global round. Otherwise, the attacks can be launched at any given global round. Figure 2 presents a comparison of the accuracy of the global model under different data poisoning attacks and with and without the proposed framework for ECG classification. Here, GM is a global model with our proposed framework under different data poisoning attacks, LS represents the global model under a label swapping attack without our proposed framework, RLF represents the global model under a random label and feature poisoning attack without our proposed framework, RL represents the global model under a random label attack without our proposed framework, and FP represents the global model under a feature poisoning attack without our proposed framework, all in a federated setting. It can be seen that the performance of the global model deteriorates significantly under the data poisoning attacks. It can also be seen that the performance of the global model was not affected by the poisoning attacks when we applied our proposed framework that could detect and eliminate almost all poisoned updates from the malicious edge before global aggregation. As mentioned previously, even if the accuracy of AM for benign and malicious sample detection is not 100%, it can accurately detect all the malicious updates because the poisoning rate is high. Hence, its malicious updates detection is still 100%. Similarly, Figure 3 presents the performance of the global model compared under four different data poisoning attacks with and without adopting the proposed framework for HAR. Since all the malicious updates are detected and removed before aggregation, the accuracy of the global model remains almost similar for a given dataset.
@@
Moreover, Figure 4 compares the global model’s accuracy under four different model poisoning attacks with and without our proposed framework for ECG classification. Here, GM shows the performance of the global model with our proposed framework, SF shows the performance of the global model under a sign flip attack without our proposed framework, SV shows the performance of the global model under a same value attack without our proposed framework, AGA shows the performance of the global model under a performance additive Gaussian noise attack without a proposed framework, and GA shows the performance of the global model under a gradient ascent attack without our proposed framework, all in a federated setting. Similarly, Figure 5 compares the global model’s accuracy under the four different model poisoning attacks with and without our proposed framework for HAR.